OpenStudy (anonymous):

Help! i give medals Find the height and surface area of the right cylinder with the height equal to the diameter (volume is 128pi cubed)

OpenStudy (anonymous):



OpenStudy (espex):

Well you know the volume is \(\pi r^2h\) you can use the dimensions you are given to solve for 'r'. Your surface area is \(2\pi rh + 2\pi r^2\) plug and chug.

OpenStudy (anonymous):

so is it \[128\pi=pir2r\]

OpenStudy (jim766):

\[128\pi^3=\pi(r^2)(2r)\]

OpenStudy (jim766):

\[128\pi^3= \pi(2)r^3\]

OpenStudy (jim766):

\[\frac{ 128\pi^3 }{ \pi }=\frac{ 2\pi(r^2) }{ \pi} \]

OpenStudy (jim766):

128pi^2=2r^2 64pi^2=r^2

OpenStudy (jim766):

8pi = r

OpenStudy (anonymous):

r=8

OpenStudy (jim766):

8pi

OpenStudy (jim766):

good?

OpenStudy (anonymous):

thank you for your help much appreciated

OpenStudy (jim766):

yw